# Learning to flock in open space by avoiding collisions and staying together

### v1

@@ Abstract
We further characterize the internal structure and dynamics of the resulting groups using liquid-state metrics and neighbour exchange rates, finding qualitative agreement with empirical observations in starling flocks. These results suggest that flocking may emerge in groups of moving agents as an adaptive response to the biological imperatives of staying together while avoiding collisions.
1 Introduction
@@
In this work, we tackle this problem through the lens of Multi Agent Reinforcement Learning (MARL) [14] considering both a centralized learning scenario, in which all particles share their training, and a concurrent one where each agent learns separately. This approach was explored in [15] where it was demonstrated that velocity alignment emerges as an effective strategy to favor cohesion. These findings were corroborated by a theoretical study of mean-field optimal stochastic control in [16]. Here we extend the scope of these work by considering a richer model for the agents’ behavior and their interactions. Our SPPs move in a two dimensional open space and only perceive the positions and orientations of a local set of topological neighbours [17, 7]. They judge their current state in terms of the mean distance to these neighbours, and take actions to modify their instantaneous direction of motion, combining alignment with the neighbours orientation and attraction towards their position. For every state, they are trained to find the optimal combination of alignment and attraction according to a cost function based on the distance to local neighbours. Regardless of the training being centralized or concurrent, we discover that cost functions that simply penalize large distances lead to a swarming behavior dominated by attraction, while the addition of a further penalty at small distances favor strong alignment at small to moderate distances, leading to a clear flocking and cohesive collective behavior similar to the one observed in starling flocks.
@@ 2.1 Dynamics
where $\textbf{r}_{i}^{t}$ is the position of the i-th agent at time t, $\textbf{s}_{i}^{t}$ its unit heading direction and the speed of the agents is kept constant and equal to $v_{0}$. In the following, without loss of generality, we put $\Delta t=1$. Agents only perceive their local environment represented by their topological (i.e. Voronoi) neighbours, the set ${V}_{i}$, and integrate their information into the local average heading direction
@@
$\textbf{R}_{i}^{t}=\Theta\left[\sum_{j\in\mathcal{V}_{i}}(\textbf{r}_{j}^{t}-\textbf{r}_{i}^{t})\right]\,.$ (3)
@@
These vectors are normalized to unity by the normalization operator $\Theta[{\bf v}]\equiv{\bf v}/|{\bf v}|$. At each time step, agents change their direction of motion, aligning towards a combination of the two vectors above, weighted by a (learnable and state-dependent) parameter $\beta\in[0,1]$, with
@@
$\langle\xi_{i}\cdot\xi_{j}\rangle=\delta_{ij}\,.$ (5)
Note that agents first update their heading via Eq. (4) and then their position by Eq. (1), a sketch of the dynamics is given in Fig. 1a. Note also that for $\beta=0$ this rule reduces to the standard Vicsek model with topological interactions[7]. The relative importance of the attraction towards the local center of mass w.r.t. the alignment interaction is measured by the action parameter $\beta$, which may depend on the local state of individual agents and will be optimized via a reinforcement learning algorithm. Note that attractive forces are expressed as an (overdamped) torque acting on the particles heading as in Ref. [8], so that particles heading always coincides with their direction of motion, a framework suitable for active agents that interact via social forces without physical collisions.
2.2 MARL Framework
Reinforcement Learning (RL) is a machine learning paradigm in which an agent learns to make decisions by interacting with the environment [14]. The agent perceives some form of sensory input from the environment – i.e. a representation of its current state – and, based on that observable, performs an action and changes its state. After it has evolved its state, it receives feedback in response to its old state, the action taken and the new state it finds itself into, according to a specific reward (or, equivalently, a cost) function which encodes the agent’s overall goals. This feedback, in turn, may reinforce or penalize the likelihood of selecting the same action again when encountering the same state. Eventually, repeating this cycle and through trial-and-error, an agent should learn an optimal policy, which maps states to actions, maximizing the cumulative reward (or, equivalently, minimizing the cost) over time. In multi-agent RL, different agents interact between themselves and with the environment. See Refs. [21, 15, 18, 19, 20] for previous applications to active matter dynamics.
@@ 2.2 MARL Framework
We store the cumulative rewards of each couple of states and actions in the so-called $Q$ matrix. States are encoded in the $Q$ matrix via the mean distance of agent $i$ to its local neighbours,
@@
where the sum is over the set $\mathcal{V}_{i}$ of Voronoi neighbours of the i-th agent and $m_{i}$ is the cardinality of such set. Actions amount to one step of the dynamics (4)-(1) with a specific choice of the parameter $\beta$ in Eq. (4) which determines the balance between aligning and attractive interactions between agents: $\beta=0$ results in a purely aligning dynamics (à la Vicsek), and $\beta=1$ gives complete attraction towards the center of mass of local neighbours, that is an attraction dominated dynamics. For the sake of simplicity, we discretize both the states and the actions – respectively with steps $\Delta d$ and $\Delta\beta$ in order to have a finite number of states $n_{s}$ and a finite number of possible actions $n_{a}$, avoiding the unnecessary computational complexity associated with continuous states and actions. State-action pairs are thus the discrete set $(d_{k},\beta_{j})$ with $k=1,\ldots,n_{s}$ and $j=1,\ldots,n_{a}$.
[…]
$\epsilon_{i}(d_{k},\beta_{j})=(1+\Omega_{i}(d_{k},\beta_{j})/\Omega_{0})^{-\nu},$ (7)
where the frequency matrix $\Omega_{i}(d_{k},\beta_{j})$ is the number of times the agent visited that particular state-action pair during the entire learning procedure, $\Omega_{0}=100$ a frequency normalization and $\nu\in(0,1)$ [22] the so-called exploration exponent.
The goal we impose to our agent group is to maintain cohesion while, at the same time, avoiding collisions between agents. Therefore, in order to evaluate the performance of the state-action pair, we adopted a cost function in the form of gentler version of a Lennard-Jones-like potential:
$\mathcal{L}(d)=\frac{a}{d}-\frac{b}{\sqrt{d}},$ (8)
@@
The goal of the learning agents is to find a policy that minimizes the cost. To do so, in the DT paradigm the $i$-th agent learns its own matrix $Q_{i}^{t}$. At each time step, only the element corresponding to the chosen state-action couple $(\tilde{d}_{i},\beta_{i})$ is updated according to the following update rule:
$Q_{i}^{t+1}(\tilde{d}_{i},\beta_{i})=Q_{i}^{t}(\tilde{d}_{i},\beta_{i})+\alpha(\tilde{d}_{i},\beta_{i})[c_{i}^{t+1}-Q_{i}^{t}(\tilde{d}_{i},\beta_{i})],$ (10)
@@
where we initialize the $Q_{i}$ matrices to 0 and the learning rate $\alpha_{i}$ decays as training proceeds according to
$\alpha_{i}(d_{k},\beta_{j})=\alpha_{0}\,(1+\Omega_{i}(d_{k},\beta_{j})/\Omega_{0})^{-\omega}.$ (11)
with $\alpha_{0}=0.005$ and $\omega\in(0.5,1)$ [22] the so-called learning exponent. In practice, as the training proceeds, the matrices $Q_{i}$ converge towards a stationary state with the optimal policy for each state $d$ defined as
$\beta^{*}(d)=\operatorname*{arg\,min}_{\beta}Q(d,\beta)\,.$ (12)
Note that in the optimization process we take into account only the immediate reward. This corresponds to a single-step Q-learning scheme, where agents only optimize short-term improvements in local cost.
While DT assumes that each agent learns by itself, in CT agents learn by shared experience. So, instead of having $N$ different $Q_{i}$ matrices and frequency matrices $\Omega_{i}$, agents instantaneously share with all the other agents the feedback they received for picking a given state-action couple, updating the same Q-matrix. Obviously, we also have a single exploration rate $\epsilon$, learning rate $\alpha$ and frequency matrix $\Omega$, albeit with a size dependent frequency normalization $\Omega_{0}=N$.
3 Results
@@
Learning is performed over $n_{e}$ training episodes, each of a duration of $T=20N$ time steps. We discretize actions and states with $\Delta\beta=1/(n_{a}-1)$ and $n_{a}=11$ actions and $\Delta d=0.2$ with $n_{d}=20$ different possible states (the 20th state also encodes for mean neighbours distances $d>20\Delta d=4$). Unless otherwise stated, we also choose the exploration and learning exponents $\nu=\omega=0.7$, but we have verified that our results are generically robust with respect to reasonable changes in all parameter choices. At the beginning of each training episode the agents’ positions and headings are randomly reinitialized (for the positions we uniformly distribute the agents in a circle with a radius $R_{0}=10\sqrt{2}$) but of course they keep track of what they learned so far (the Q matrix is not reinitialized). Unless differently stated, we also fix $v_{0}=0.2$ and $\eta=0.3$ in the agents dynamics.
@@
Figure 2: MDCM and OP for both CT (blue, $n_{e}=300$) and DT (red, $n_{e}=1000$). (a) MDCM in the last training episode for a group of $N=100$ agents. (b) OP in the last training episode ($N=100$). Insets in (a) and (b) are averages over single episodes. (c) Finite size scaling of the average OP. Inset: Finite size scaling of the average MDCM, the dashed black line marks the power law $\sim\sqrt{N}$. (d) OP vs. noise amplitude for $N=100$. Open red symbols refer to DT with $\nu=\omega=0.7$, while full red symbols to DT with $\nu=0.5$ and $\omega=0.99$.
3.1 Onset of cohesive flocking
@@
We have verified numerically that our RL algorithm leads to a state of cohesive flocking in a relatively small number of training episodes, both in the CT and DT paradigms. In particular, flocking alignment is quantified as usual via the scalar polar order parameter (OP)
@@
For an ordered ensemble of mutually correlated agent headings $\Phi(t)$ is finite and of order one, while it tends to zero for uncorrelated and disordered headings. In finite flocks, in the latter case one has indeed $\Phi(t)\sim 1/\sqrt{N}$ due to finite fluctuations. Furthermore, group cohesion can be assessed via the mean distance to the group center of mass (MDCM)
@@
We established the scaling of our training algorithm with the number of agents testing groups up to $N=1600$ agents (a typical size for starling flocks observed in the field [17]). As shown in Fig. 2c, the average OP at the end of training for CT converges towards a finite asymptotic value $\Phi\approx 0.94$, while the flock size $\Gamma$ grows like $\sqrt{N}$, signaling that the flock density remains constant as the group size is varied. DT, on the other hand, seems to be less robust, with $\Phi\approx 0.84$ and a large flock size $\Gamma$ signaling a partial loss of cohesion at $N=1600$. This may be partially mitigated by a more careful choice of the exploration ($\nu=0.5$) and learning ($\omega=0.99$) exponents, a choice that increases the late time exploration rate and reduces the learning rate. A more careful analysis of finite size scaling in the DT regime is left for future work.
@@
Finally, in Appendix A we characterize the structure of cohesive flocks, showing that our flocks behave as a rather structureless nonequilibrium liquid, with a finite neighbours exchange rate [23] and a pair distribution function qualitatively matching experimental observations in starling flocks [24].
3.2 Analysis of the Q-matrix
The (normalized) final $Q$-matrices for CT and DT are shown (respectively) in Fig. 3a and Fig. 3c for $N=100$. They are characterized by two main regimes, a first one for states $d>d^{*}\approx 2$ and a second one at shorter distances, $d<d^{*}$. At large mean distances to neighbours, the optimal actions for the agents, given by Eq. (12), involve a combination of alignment with, and attraction to, their neighbours, with $\beta$ seemingly randomly scattered in the interval $(0,1]$. In Appendix B we further show that the exact value of the optimal action $\beta$ for $d>d^{*}$ is not significant, and that different (random) combinations of alignment and attraction lead to essentially the same cohesive flocking behavior.
Note finally the exception of the first state, $d<\Delta d=0.2$, where attraction seems to be the optimal action in both CT and DT. This rather puzzling result is a consequence of the discrete nature of our dynamics: when $d<\bar{d}$ the cost function rewards actions that increase the mean distance to neighbours $d$ and penalizes those that further reduce it. For $d>v_{0}$ this simply rules out attraction, resulting in agents with commonly aligned headings that separate slowly due to noisy fluctuations in their individual headings. However, when extremely close, $d\leq v_{0}$, strong attraction may lead to large differences in neighbours’ heading, so that nearby agents may easily pass each other and find themselves, after a single time-step, separated by a larger distance. In practice, at distances smaller than $v_{0}$, large $\beta$ may act like an effective repulsion and be favored by the diverging cost function. We have indeed verified that, for a reduced speed $v_{0}\ll\Delta d$ (in particular, we have chosen $v_{0}=0.05$), this effect is largely mitigated, with a typical optimal policy for the smallest mean neighbour distance state ($d<\Delta d$) dominated by alignment, $\beta\in[0,0.2]$. In Appendix B we provide further evidence that this effect is due to discretization of the dynamics and show that forcing $\beta=0$ for $d<v_{0}$ still leads to cohesive flocking.
3.3 A swarming counterexample
We conclude by discussing the role of the cost function (8) divergence at short distances in the onset of flocking behavior. Such a divergence can be removed by considering instead $\mathcal{L}(d+z)$ with a positive shift $z$. In particular, with the choice $z=\bar{d}=1$ the minimum of the cost function is shifted to $d=0$, completely removing penalties for arbitrarily small agent distances $d$. We have verified, as shown in Fig. 4, that this choice suppresses flocking and only leads to cohesive swarming. The analysis of the trained $Q$-matrix indeed shows that in this case the orientational dynamics is dominated by attraction ($\beta\approx 1$).
3.4 Boundary agents and Voronoi based cost
Since Voronoi cells at the group boundary have, by definition, an infinite area, corresponding to a vanishing cost function $\mathcal{L}$, our new choice highlights the cost/reward difference between agents located inside and those at the boundary of the flock. Therefore, actions resulting in agents moving from the boundary to the interior of the group will be strongly rewarded, while moving from the interior to the border is strongly penalized. Moreover, agents at the boundary get zero cost/reward for any action that does not move them to the inside of the flock, regardless of any change in the distance to their Voronoi neighbours. We have numerically tested this new learning scheme both for centralized and decentralized training, but we have not found any qualitative difference with cost/rewards based on the distance to neighbours. Agents still learn similar optimal policies leading to cohesive flocking, only with a slightly larger typical flock size, a consequence of the absence of reward for boundary agents that get closer to their neighbours.
4 Discussion
Interestingly, the learned flocking dynamics shows characteristic structural features, such as a slow mixing rate and a structureless pair distribution function, reminiscent of natural bird flocks, reinforcing the biological plausibility of the learned behavior. While our results have been obtained in two spatial dimensions, one could argue that – due to gravity – natural birds also fly predominantly in a quasi-2D plane [28, 29].
We finally note that our results may also prove useful beyond animal behavior, particularly in the field of swarm robotics [30]. In the future, it may be interesting to generalize these results to mixed-agent populations and environments with external threats.
Acknowledgments
Appendices
Appendix A Flock Structure

### v2

@@ Abstract
Active matter, Reinforcement learning
I Introduction
@@
In this work, we tackle this problem through the lens of Multi Agent Reinforcement Learning (MARL) [4] considering both a centralized learning scenario, in which all particles share their training, and a concurrent one where each agent learns separately. This approach was explored in [16] where it was demonstrated that velocity alignment emerges as an effective strategy to favor cohesion. These findings were corroborated by a theoretical study of mean-field optimal stochastic control in [5]. Here we extend the scope of these work by considering a richer model for the agents’ behavior and their interactions. Our SPPs move in a two dimensional open space and only perceive the positions and orientations of a local set of spatially balanced topological neighbours [3, 20, 8]. They judge their current state in terms of the mean distance to these neighbours, and take actions to modify their instantaneous direction of motion, combining alignment with the neighbours orientation and attraction towards their position. For every state, they are trained to find the optimal combination of alignment and attraction according to a cost function based on the distance to local neighbours. Regardless of the training being centralized or concurrent, we discover that cost functions that simply penalize large distances lead to a swarming behavior dominated by attraction, while the addition of a further penalty at small distances favor strong alignment at small to moderate distances, leading to a clear flocking and cohesive collective behavior similar to the one observed in starling flocks.
@@ II.1 Dynamics
where $\textbf{r}_{i}^{t}$ is the position of the i-th agent at time t, $\textbf{s}_{i}^{t}$ its unit heading direction and the speed of the agents is kept constant and equal to $v_{0}$. In the following, without loss of generality, we put $\Delta t=1$. Agents only perceive their local environment represented by their topological (i.e. Voronoi) neighbours, the set ${V}_{i}$, a spatially balanced version of metric free interactions that grants optimal stability [8]. They integrate this information into the local average heading direction
@@
$\textbf{R}_{i}^{t}=\Theta\left[\sum_{j\in\mathcal{V}_{i}}(\textbf{r}_{j}^{t}-\textbf{r}_{i}^{t})\right]\,$, (3)
@@
where the sum is over the set $\mathcal{V}_{i}$ of Voronoi neighbours of the i-th agent. These vectors are normalized to unity by the normalization operator $\Theta[{\bf v}]\equiv{\bf v}/|{\bf v}|$. At each time step, agents change their direction of motion, aligning towards a combination of the two vectors above, weighted by a (learnable and state-dependent) relative coupling strength parameter $\beta\in[0,1]$, with
@@
$\langle\xi_{i}^{t}\cdot\xi_{j}^{t^{\prime}}\rangle=\delta_{ij}\delta_{t,t^{\prime}}\,.$ (5)
Agents first update their heading via Eq. (4) and then their position by Eq. (1), a sketch of the dynamics is given in Fig. 1a. Note that for $\beta=0$ this rule reduces to the standard Vicsek model with topological interactions [20], while $\beta=1$ gives complete attraction towards the center of mass of local neighbours, that is an attraction dominated dynamics. The relative coupling strength is thus the crucial parameter that will be optimized via the reinforcement learning algorithm described in the next subsection. Note finally that attractive forces are expressed as an (overdamped) torque acting on the particles heading as in Ref. [23], so that particles heading always coincides with their direction of motion, a framework suitable for active agents that interact via social forces without physical collisions.
II.2 MARL Framework
[…]
@@ II.2 MARL Framework
The raw observation of each agent $i$ is given by the relative positions and headings of its Voronoi neighbours, while its policy observation (i.e. the information on which the policy is conditioned) is given by the mean distance of agent $i$ to its local neighbours,
@@
where $m_{i}$ is the cardinality of the set $\mathcal{V}_{i}$ of Voronoi neighbours. Actions amount to the choice of a specific value for the relative coupling strength $\beta$ with which the agent performs one step of the dynamics (4)-(1).
With a slight abuse of terminology, in the following we will use the term “state” for the policy observation $d(i)$ (not to be confused with the microscopic agent state $({\bf r}_{i},{\bf s}_{i})$, and denote the chosen $\beta$ with the term “action”. This is a standard use in RL where one speaks generically of state/action pairs. For the sake of simplicity, we also discretize both the states and the actions in order to have a finite number of states $n_{s}$ and a finite number of possible actions $n_{a}$, avoiding the unnecessary computational complexity associated with continuous states and actions (we describe the details of the discretization in the Results section) .
State-action pairs are thus a discrete set of pairs $(d,\beta)$ defining the entries of the so-called $Q$-matrix that in RL represent the estimate for the expected cost for each state-action pair and thus determines the optimal policy
$\beta^{*}(d)=\operatorname*{arg\,min}_{\beta}Q(d,\beta)\,.$ (7)
We first discuss how the cost is encoded in the Q matrix and afterwards how actions are selected during learning. The goal we impose to our agent group is to maintain cohesion while, at the same time, avoiding collision risks between agents (that is relative distances getting too small). Therefore, in order to evaluate the performance of the state-action pair, we adopted a cost function in the form of a gentler version of a Lennard-Jones-like potential:
$\mathcal{L}(d)=\frac{a}{d}-\frac{b}{\sqrt{d}}$ (8)
@@
First consider the decentralized training (DT) paradigm, where each agent is characterized by a different $Q_{i}$ matrix. At each time-step $t$, agent $i$ chooses an action $\beta_{i}$ according to its state $d_{i}$, evolves its position and heading via Eqs. (1)-(4), and computes the corresponding cost (9). The $Q_{i}$-matrix element corresponding to a chosen state-action couple $(d,\beta)$ is updated according to the following update rule
$Q_{i}^{t+1}(d,\beta)=Q_{i}^{t}(d,\beta)+\alpha(d,\beta)[c_{i}^{t+1}-Q_{i}^{t}(d,\beta)],$ (10)
@@
where we initialize the $Q_{i}^{0}$ matrices to 0 and the learning rate $\alpha_{i}$ decays as training proceeds according to
$\alpha_{i}(d,\beta)=\frac{\alpha_{0}}{(1+\Omega_{i}(d,\beta)/\Omega_{0})^{\omega}},$ (11)
with $\alpha_{0}\ll 1$ and $\omega\in(0.5,1)$ [35] the so-called learning exponent. The frequency matrix $\Omega_{i}(d,\beta)$ is the number of times the agent visited that particular state-action pair during the entire learning procedure and $\Omega_{0}$ is a frequency normalization. When an agent selects a state-action couple $(d,\beta)$ the information it will learn will be weighted less the more often that state-action couple has been visited. This ensures the convergence of the $Q$-matrix towards a stationary state.
The goal of RL agents is to minimize the cumulative cost they collect over time, and typically this may be done by choosing the current optimal action via Eq. (7). On the other hand, agents need to explore all the possible combinations of states and actions in order to find the real optimum. The exploration-exploitation tradeoff is a central topic in RL algorithms and, in order to be correctly addressed, one needs to balance the two paradigms correctly. For this reason we decided to induce exploration of new state-action pairs using the so called $\epsilon-greedy$ exploration scheme. At each time step, agents choose the current optimal action (7) with probability $p=1-\epsilon_{i}(d,\beta)$ or, with complementary probability $\epsilon_{i}(d,\beta)$ a random action $\beta$ chosen with equal probability between the $n_{a}$ different options. The exploration rate $\epsilon_{i}$ is a function of state-actions pairs and decreases during learning according to the decay rule
$\epsilon_{i}(d,\beta)=(1+\Omega_{i}(d,\beta)/\Omega_{0})^{-\nu},$ (12)
with an exploration exponent $\nu\in(0,1)$ [35]. Random exploration avoids the training dynamics to be stuck in sub-optimal minima; analogously to the learning rate, the exploration rates decays towards zero as state action pairs are repeatedly visited, ensuring convergence of the learning algorithm.
Learning is typically organized in $n_{e}$ training episodes of $T$ time-steps (in a time steps all agent evolve once with a synchronous dynamics), with the initial positions and headings of the agents reinitialized at the beginning of each episode. The $Q$-matrices, learning rates and exploration rates, of course, are not reinitialized every episode. At the end of learning, one expects the $Q$-matrices to have practically converged towards a stationary state and one can define a learned dynamics where agents always select the optimal action for each given state via Eq. (7).
Centralizing training (CT) follows the same procedure, but in this case agents share the same information during learning, with single $Q$ and frequency matrices, $Q_{i}=Q$, $\Omega_{i}=\Omega$ and learning and exploration rates, $\alpha_{i}=\alpha$, $\epsilon_{i}=\epsilon\,\,,\,\forall i$.
III Results
Learning is performed over $n_{e}$ training episodes, each of a duration of $T=20N$ time steps. We discretize actions and states with $\Delta\beta=1/(n_{a}-1)$ and $n_{a}=11$ actions and $\Delta d=0.2$ with $n_{d}=20$ different possible states (the 20th state also encodes for mean neighbours distances $d>20\Delta d=4$). We have verified that our results are robust with respect to reasonable increases in $n_{d}$ and $n_{a}$. For what concerns the learning parameters introduced in the previous section, we have performed a preliminary exploration of different values in the convergence windows of RL theory [35] in order to adapt them to our specific problem. While we cannot claim optimality in a rigorous sense, we have settled for $\alpha_{0}=0.005$, $\Omega_{0}=N$ and $\nu=\omega=0.7$ for centralized training. These choices seem to grant a sufficiently fast convergence towards a stationary state, with fairly robust results w.r.t. small changes in these parameters. As we will show in Sec. III.1, decentralized training is more delicate. For sufficiently small flocks, $N<800$, these parameters, with the frequency normalization $\Omega_{0}=100$ are still efficient, while for larger flocks we obtained better cohesion with $\nu=0.5$ and $\omega=0.99$.
@@ III Results
At the beginning of each training episode the agents’ positions and headings are randomly reinitialized (for the positions we uniformly distribute the agents in a circle with a radius $R_{0}=10\sqrt{2}$) but of course they keep track of what they learned so far (the Q matrix is not reinitialized). Unless differently stated, we also fix $v_{0}=0.2$ and $\eta=0.3$ in the agents dynamics.
@@
Figure 2: Mean distance to the centre of mass (MDCM) and order parameter (OP) for both centralized (CT) (blue, $n_{e}=300$) and decentralized (DT) training (red, $n_{e}=1000$). (a) MDCM in the last training episode for a group of $N=100$ agents. (b) OP in the last training episode ($N=100$). Insets in (a) and (b) are averages over single episodes. (c) Finite size scaling of the average OP. Inset: Finite size scaling of the average MDCM, the dashed black line marks the power law $\sim\sqrt{N}$. (d) OP vs. noise amplitude for $N=100$. In panels (c)-(d) open red symbols refer to DT with $\nu=\omega=0.7$, while full red symbols to DT with $\nu=0.5$ and $\omega=0.99$.
III.1 Onset of cohesive flocking
@@
We have verified numerically that our RL algorithm leads to a state of cohesive flocking in a relatively small number of training episodes, both in the CT and DT paradigms, see for instance Video1 in Supplementary. In particular, flocking alignment is quantified as usual via the scalar polar order parameter (OP)
@@
For an ordered ensemble of mutually correlated agent headings $\Phi(t)$ is finite and of order one, while it tends to zero for uncorrelated and disordered headings. In finite flocks, in the latter case one has indeed $\Phi(t)\sim 1/\sqrt{N}$ due to finite fluctuations. Furthermore, group cohesion can be assessed via the mean distance to the group center of mass (MDCM) which estimates the typical flock size
@@
We established the scaling of our training algorithm with the number of agents testing groups up to $N=1600$ agents (a typical size for starling flocks observed in the field [9]). As shown in Fig. 2c, the average OP at the end of training for CT converges towards a finite asymptotic value $\Phi\approx 0.94$, while the flock size $\Gamma$ grows like $\sqrt{N}$, signaling that the flock density remains constant as the group size is varied. Decentralized training, on the other hand, seems to be less robust, with $\Phi\approx 0.84$ and a large flock size $\Gamma$ signaling a partial loss of cohesion at $N=1600$. This is due to the increasing difficulties of individual agents to explore the full range of states $d_{i}$ in a finite time when $N$ is increased. In practice, some agents do not learn an optimal policy in a reasonable training time. This difficulty may be partially mitigated by a change of the exploration ($\nu=0.5$) and learning ($\omega=0.99$) exponents, a choice that increases the late time exploration rate at the expense of a faster decreasing learning rate. As shown in Fig. 2, these values lead to nearly identical results for $N<800$ but result in better cohesion at larger system sizes. A more careful analysis of the finite size scaling of training in the DT regime is left for future work.
@@
III.2 Flock Structure
We characterized the structure of cohesive flocks, showing that our flocks behave as a rather structureless nonequilibrium liquid, with a finite neighbours exchange rate [30] and a pair distribution function qualitatively matching experimental observations in starling flocks [2].